OpenStudy (90890908967777878888):

i need help because i'm kind confused may some1 help plz srry i'll reward u with a medal and because a fan 20 f+10 A.f(2+10) B.2(f+5) C.2(5+10)

OpenStudy (alexandervonhumboldt2):

WHAT DO YOU NEED TO DO? MORE INFO PLEASE

OpenStudy (90890908967777878888):

i need to use the Distributive Property to write an equivalent expression.

OpenStudy (trantom):

what does "20 f+10" mean? 20*(f+10) or 20f + 10?

OpenStudy (alexandervonhumboldt2):

there is no equaling expression

OpenStudy (alexandervonhumboldt2):

A.f(2+10)=2f+20 B.2(f+5)=2f+10 C.2(5+10)=10+20=30

OpenStudy (90890908967777878888):

hold on here are the question to the hold test

OpenStudy (90890908967777878888):

1. 2f + 10 (1 point) f(2 + 10) 2(f + 5) 2(f + 10) 2. 20 + 24g (1 point) 4(5 + 6g) 20(1 + 4g) 5(4 + 6g) 3. 32b – 20 (1 point) 4(8b – 20) 2(32b – 10) 4(8b – 5) For questions 4–5, simplify the expression. 4. 24h + 3d – 10h + 8d (1 point) 14h + 11d 14h – 5d 25hd 5. 12s + 4 + 4s – 2 (1 point) 18s 2(8s + 3) 2(8s + 1)*

OpenStudy (90890908967777878888):

@AlexandervonHumboldt2

OpenStudy (trantom):

Well here, the distributive property works like this: 4(f+2) 4f + 4(2) 4f + 8 so in this example, 4(f+2) = 4f + 8

OpenStudy (90890908967777878888):

okay do u have the answers

OpenStudy (trantom):

yes, but the point is for you to get the answers If you don't try and always copy people's answers, you're not going to learn, and then when you take the SAT and math problems you don't know are on it you won't know how to do them which might prevent you from getting into a good college, which will have consequences lasting your whole life If you learn how to do it, then you won't have this problem, and math will actually be easier for you So look at my example I showed you and apply it to problem 1 I'll go through it with you but I'm not going to give you the answers right away, ok?